) Given that 3 + i is a root of the equation z − 3z − 8z + 30 = 0, find the
remaining roots.
1
Expert's answer
2020-05-20T18:37:29-0400

Solution: According to the complex conjugate root theorem: if 3+i is a root, then 3-i is a root too. In this way the original polynomial P=z3-3z2-8z+30=(z-z1)(z-3-i)(z-3+i)=(z-z1)((z-3)2-i2)=(z-z1)(z2-6z+10);To find remaing roots - divide the original polynomial by z2-6z+10:z3-3z2-8z+30=z*(z2-6z+10)+3*(z2-6z+10)=(z+3)(z2-6z+10);I.e. the remaining root is equal to -3.
Answer: 3-i, -3.